OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (tkhunny):

Unfortunately, with only a small number of values, the "suggested pattern" doesn't actually exist.

OpenStudy (anonymous):

I have these choices though @tkhunny

OpenStudy (anonymous):

oops, actually these are my choices

OpenStudy (paxpolaris):

lol better

OpenStudy (anonymous):

a=-9 d=-3-(-9)=-3+9=6 l=81 l=a+(n-1)d 81=-9+(n-1)6 n=? then write as A.P. series

OpenStudy (perl):

thats a sum of a sequence that is linear

OpenStudy (perl):

plug in the top index, 15 -9 + 6*15 , that will give you the last value

OpenStudy (tkhunny):

Well, it's not \(\infty\). I wouldn't even look at "54", either.

OpenStudy (anonymous):

so a? @perl @surjithayer

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

thank you so much! could you help me with a couple other different questions?

OpenStudy (tkhunny):

Why are you tagging anyone? You have enough information to solve it. Write down the correct answer and move onl Post new question on new threads and ALWAYS show YOUR work. Always! No excuses.

OpenStudy (anonymous):

81+9=(n-1)6 n-1=81/6=15 n=15+1=16 n from 0 to 15 gives 16 terms so first option matches this

OpenStudy (anonymous):

@surjithayer thank you(:

OpenStudy (anonymous):

yw